# MNSC-0001-2024.00

### v1

@@ 6 Theoretically Evaluation
Proof: The IBFT mechanism we have introduced establishes a clear association between the AIGC product and the transaction ID provided by Producer and Provider. Since forged $Txids^{\prime}$ cannot pass IBFT.check(), preventing them from being mapped within BF, thereby confirming their irrelevance to the AIGC product. This mechanism effectively hinders attackers from posting fraudulent transactions on the blockchain to interfere with the accurate determination of copyright ownership for AIGC products.
@@
Proof: Auditors responsible for copyright management are composed of copyright certification institutions, courts, and other authoritative bodies, which also serve as nodes within the AIGC-Chain network. They rely on dependable, immutable data from the blockchain. The decentralized nature of the blockchain necessitates consensus among all nodes for audit conclusions, significantly reducing the risk of malicious auditor actions and ensuring the fairness and security of copyright management. Moreover, as nodes on the blockchain, Auditors risk having their node weights diminished after making mistakes, which can adversely affect their interests. Consequently, auditors are incentivized to avoid bias and ensure that their decisions are accurate, fair, and trustworthy.
@@ 7.1 Implementation Overview
AIGC-Chain is compatible for deployment across various blockchain platforms. Recognizing the lack of a gas-like evaluation system in Hyperledger Fabric for smart contract performance assessment, we opted to establish a private Ethereum network to quantify the performance of AIGC-Chain. The blockchain is implemented in Go and utilizes JavaScript with Node.js for smart contracts. In our test network, we have established an order node along with two peer nodes. Each node is hosted within its own Docker container to integrate into the system. All nodes operate on virtual machines endowed with 4GB of RAM and powered by the Ubuntu 16.04 operating system. Besides, the size of the used IBF was set to $m$=10000, and the false positive rate was set at 0.01.
@@ 7.2 Blockchain Evaluation
Figure 8: Transaction Cost Analysis for AIGC-Chain
Fig. 8 reveals the transaction costs associated with different procedures within the AIGC-CHain. The gas cost for $TX_{req}$ is particularly high, reaching up to 217,370 gas, which is mainly attributed to the extensive write operations necessary for the initialization of an AIGC product on the blockchain. Conversely, the stage of Copyright Management does not incur any transaction costs since no data needs to be written onto the blockchain in this step.
Figure 9: Execution Cost Analysis for AIGC-Chain
The execution gas cost represents the computational resource expenditure required for the execution of smart contracts, and it also encompasses the associated costs for retrieving data from the blockchain. On the AIGC-Chain, the distribution of execution gas costs across five steps is depicted in Fig. 9. The initialization in $TX_{req}$ incurs a higher execution cost, whereas the gas expenses for the other stages are comparatively lower. Furthermore, compared to Fig. 8, the cost of execution gas is significantly lower than that of transaction gas.
7.3 IBFT Evaluation
We perform a comparative analysis of the proposed IBFT algorithm with Fast Query (Chen et al. 2023c). IBFT achieves rapid transaction verification by utilizing a designated set within the world state, whereas Fash Query accelerates search operations through an array that specifically contains copyright transactions associated with product P. In contrast, the traditional method necessitates a full scan of the stored information on the blockchain for query resolution. The time complexity for query execution among these methods is presented in Table 2, with m representing the total number of transactions on the blockchain and n indicating the count of transactions pertinent to product P.
Table 2: COMPARISON OF TIME COMPLEXITY
@@ 7.3 IBFT Evaluation
IBFT O(1)
In each test iteration, we generate 1000 transactions on the blockchain, progressively adjusting the proportion of transactions related to product P. Assuming the transaction to be queried pertains is the initial transaction of this product, the time cost for querying the validity of this transaction is illustrated in Fig. LABEL:test1. Normal Query invokes smart contracts to query all the transactions on the blockchain, resulting in a fixed yet higher time cost. Fast Query, however, only traverses transactions related to P, leading to a negligible increase in time cost as the proportion of P-related transactions on the blockchain increases. In the case of IBFT, each query accesses BF stored in the world state, necessitating only a single on-chain query and a single invocation of the query algorithm. Consequently, its time cost remains constant and is unaffected by the transaction proportion. We further conduct an experiment setting the proportion of P-related transactions at 50% and measure the associated time costs, shown in Fig. LABEL:test2. This demonstrates that IBFT facilitates efficient retrieval and validation of transaction information, substantially enhancing the auditability and validation strength of copyright-related transactions on the blockchain.
(a) The Time Cost with the Increasing Proportion of Product
(b) The Time Cost at a Fixed Proportion of Product
Figure 10: The Time Cost of IBFT
@@ 8 Conclusion
This paper introduces AIGC-Chain, a pioneering technology that integrates blockchain to create a full lifecycle copyright recording system for managing AIGC products. AIGC-Chain captures and records data throughout the lifecycle of AIGC products on the blockchain, leveraging its immutability and decentralized nature to ensure the secure storage and traceability of AIGC copyrights, thereby providing robust data support for the management of copyright ownership. Additionally, this paper introduces an efficient copyright transaction tracing method based on the Indistinguishable Bloom Filter, named IBFT, which aims to enhance the efficiency and credibility of copyright audits. This suite of innovations robustly safeguards the legal rights of creators and provides a reliable framework for proof collection and review in copyright disputes, thereby promoting the sustained growth and prosperity of the AIGC industry. As future directions, we aim to enhance the design of various aspects within the AIGC-Chain and to evaluate potential additional risks to AIGC copyrights, while also exploring appropriate countermeasures.

### v2

@@ 6 Theoretically Evaluation
Proof: CTrace establishes a clear association between AIGC product and transaction ID provided by User and Provider. Since forged $Txid^{\prime}$ cannot pass CTrace.check(), preventing them from being mapped within the blockchain, thereby confirming their irrelevance to the AIGC product. This mechanism effectively hinders attackers from posting fraudulent transactions on the blockchain to interfere with the accurate determination of copyright ownership for AIGC products.
@@
Proof: Auditors responsible for copyright management are composed of copyright certification institutions, courts, and other authoritative bodies, which also serve as nodes within the AIGC-Chain network. They rely on dependable, immutable data from the blockchain. The decentralized nature of the blockchain necessitates consensus among all nodes for audit conclusions, significantly reducing the risk of malicious auditor actions and ensuring the fairness and security of copyright management. Furthermore, as participants on the blockchain, Auditors face the potential reduction in their weight following errors, which could negatively impact their standing and interests. This creates a motivation for auditors to uphold impartiality and strive for precision, ensuring that their judgments are accurate, equitable, and reliable.
Theorem 4. Parameters $m$ = 288 and $k$ = 10 are feasible and secure for CTrace under a false positive probability of 0.1%, ensuring efficient and reliable copyright verification for AIGC products.
Proof: The security and feasibility of CTrace depend on its ability to balance false positive probability $P_{fp}$, the size of array $m$, and the number of hash functions $k$. For a given $n$ (number of elements inserted), the false positive rate is bounded by:
$P_{fp}=\left(1-e^{-\frac{kn}{m}}\right)^{k}$
To achieve $P_{fp}\leq$ 0.1%, we derive the relationship between $m$, $k$, and $n$. The optimal number of hash functions $k$ minimizes $P_{fp}$ for fixed $m$ and $n$:
$k_{\text{opt}}=\frac{m}{n}\ln 2$
For $n\leq$ 20, substituting $m$ = 288 and $k$ = 10:
$P_{fp}$ $=\left(1-e^{-\frac{10\times 20}{288}}\right)^{10}$
$=\left(1-e^{-0.694}\right)^{10}$
$\approx(1-0.500)^{10}=0.500^{10}\approx 9.76\times 10^{-4}$
This yields $P_{fp}\approx$ 0.0976% , which meets the requirement $P_{fp}\leq$ 0.1%. The parameters are therefore feasible.
For $P_{fp}$ = 0.0976%, an attacker attempting to spoof $10^{4}$ AIGC products would trigger $\sim$ 9.76 false positives on average. This low rate makes systematic exploitation statistically impractical.
The bit density $\rho$ of CTrace is:
$\rho=1-e^{-\frac{kn}{m}}=1-e^{-\frac{10\times 20}{288}}\approx 0.500$
With $\rho\approx$ 50%, the filter maintains stable performance for $n\leq$20, avoiding accelerated false positive growth.
Using $k$ = 10 hash functions introduces $O(k)$ operations for insertion/query. Modern cryptographic hash functions (e.g., SHA-3) execute these operations in $\mu$s-scale, ensuring real-time compatibility with blockchain transactions.
Therefore, the parameters $m$ = 288 and $k$ = 10 satisfy $P_{fp}\leq$ 0.1% for $n\leq$ 20, maintaining security and feasibility.
Theorem 5. An attacker cannot feasibly manipulate CTrace to corrupt copyright verification proof.
Proof: The security of CTrace relies on the immutability of its bit array and the irreversibility of hash functions. To force a false positive, an attacker must either:
  • Reverse-engineer hash outputs: Cryptographic hashes (e.g., SHA-256) are preimage-resistant, making it infeasible to find an input $x^{\prime}$ that collides with a legitimate copyright entry $x$.
  • Overwhelm the filter: Saturating the filter requires inserting $n^{\prime}$ elements such that $P_{fp}$ exceeds 0.1%. For $m$ = 288, achieving $P_{fp}>$ 0.1% requires:
$n^{\prime}$ $>\frac{m\ln(1-p^{1/k})}{-k}=\frac{288\times\ln(1-(10^{-3})^{1/10})}{-10}$
$\approx\frac{288\times(-0.693)}{-10}\approx 20.0$
Since each insertion requires blockchain consensus, surreptitiously adding ( $n^{\prime}>$ 20 ) entries is detected and prevented by CTrace. Thus, the design of CTrace is intrinsically resistant to adversarial interference, ensuring the correctness of AIGC copyright verification.
@@ 7.1 Implementation Overview
AIGC-Chain demonstrates cross-platform compatibility with mainstream blockchain platforms. To objectively evaluate system performance metrics, we constructed a private Ethereum testnet implementing a Proof-of-Authority (PoA) consensus mechanism. The blockchain core is developed in Go (Golang v1.19), while smart contracts are implemented as Solidity (v0.8.0) bytecode executed through the Ethereum Virtual Machine (EVM). Our experimental network architecture comprises 8 validator nodes and one bootnode, with each node instance containerized in isolated Docker environments (v20.10.18) to ensure resource partitioning. All network components operate on KVM-based virtual machines configured with 4GB DDR4 memory, and Ubuntu 22.04 LTS. Besides, the size of bloom filter used in CTrace was set to $m$=288, and the false positive rate was set at 0.1%.
Figure 7: Performance Analysis of AIGC-Chain
@@ 7.2 Blockchain Evaluation
Fig. 7 provides a comparative analysis of the transaction gas costs associated with various procedures within the AIGC-Chain. It is observed that the gas cost for $Tx_{req}$ reaches a maximum of 217,370 gas units, which is notably higher than the costs incurred by the subsequent procedures. This increased expenditure is due to the resource-intensive write operations required for the generation of an AIGC product on the blockchain, including the registration of metadata, the generation of cryptographic hashes, and updates to the immutable ledger. Conversely, $Tx_{verification}$ incurs no transaction costs, as it is conducted off-chain through decentralized verification protocols and does not require any on-chain data writing. The metric for execution gas cost reflects the computational resource usage during the execution of smart contracts, including both algorithmic processing and on-chain data retrieval. As depicted in Figure 7, the AIGC-Chain exhibits a varied distribution of execution gas costs across its five main workflow stages. $Tx_{req}$ accounts for the highest execution costs, attributable to the intricate contract logic that involves multi-signature authentication and decentralized storage mechanisms.
Figure 8: Execution Efficiency of CTrace
7.3 Evaluation of CTrace
This research conducts a comparative evaluation of CTrace and Fast Query[9], with a focus on their respective optimization strategies for data retrieval. CTrace facilitates rapid transaction verification by leveraging a bloom filter within the world state in a systematic manner. In contrast, Fast Query optimizes search efficiency by using a single linked list to index only the copyright transactions related to product P. Traditional methods, however, necessitate exhaustive scans of the data on the blockchain for query resolution. The time complexity for executing queries using these methods is detailed in Table II, where m denotes the total number of transactions on the blockchain and n signifies the number of transactions relevant to product P. This comparative analysis quantitatively demonstrates the superiority of CTrace in data verification.
TABLE II: COMPARISON OF TIME COMPLEXITY
@@ 7.3 Evaluation of CTrace
CTrace O(1)
In the performance evaluation phase, we conducted rigorous comparative experiments between traditional Bloom filters and CTrace. Each experimental group involved standardized sequences of 10,000 insert and query operations separately, with each experiment repeated 30 times to eliminate systemic errors through averaging. As quantitatively demonstrated in Fig. 8, CTrace exhibits measurable throughput degradation compared to conventional Bloom filters. This performance differential primarily stems from its intrinsic privacy-preserving mechanisms. While this design decision incurs additional computational costs, it successfully prevents the avenues for retroactively inferring the original data through statistical analysis techniques. Nevertheless, this performance trade-off is deemed reasonable and acceptable within the domain of privacy-enhanced data structures.
Furthermore, we compared the performance of CTrace with traditional query methods and Fast Query, which relies on a single linked list, focusing on query time and storage overhead. Fig. 9(a) illustrates the query time costs of these three methods under varying data volumes. Traditional methods require invoking smart contracts to access transaction data on the blockchain, resulting in a consistent but significant time cost. Fast Query, with its off-chain single linked list storage, experiences increasing query times with the expansion of data. In contrast, CTrace, by only accessing the bloom filter stored in the world state and performing a single hash mapping operation, maintains a constant query time regardless of storage size. Fig. 9(b) depicts the storage costs under different data volumes, highlighting that storage requirements of CTrace remain constant and do not increase with data accumulation, in contrast to Fast Query. As the amount of data increases, CTrace will significantly reduce storage overhead compared to Fast Query, demonstrating superior space efficiency.
Figure 9: Performance Analysis of CTrace
@@ 8 Conclusion
This paper confronts the copyright challenges associated with AIGC and introduces a trustworthy AIGC copyright management system for AIGC. It meticulously records intermediate data throughout the full lifecycle of AIGC and anchors it within a decentralized blockchain system, thereby establishing a robust multi-party supervision framework for the fair and reliable management of AIGC Copyrights. The paper further develops a copyright tracing method utilizing Bloom filters, named CTrace, which markedly improves the efficiency and accuracy of blockchain transaction queries and verifications, and furnishes a solid proof base for adjudicating copyright disputes. These innovations robustly safeguard the legitimate rights and offer a reliable framework for the collection and verification of proof in copyright disputes, promoting the sustainable growth and vibrancy of AIGC. As future directions, we plan to divide the copyright management of AIGC into three distinct phases: model training, requirement formulation, and content generation. We aim to delve into the copyright concerns that may emerge at each phase and to propose effective resolution strategies.